The research team would like to estimate the average gender difference in age.

Compute a 90% estimate and confidence interval for the difference in the average age of Female and Male employees.

What is the margin of error of the confidence interval ? (Use 3 decimals)

Employee Number1 Annual Salary Gender Age (years) Experience (years) Training Level
1 $     32,368 F 42 3 B
2 $     53,174 M 54 10 B
3 $     52,722 M 47 10 A
4 $     53,423 M 47 1 B
5 $     50,602 M 44 5 B
6 $     49,033 M 42 10 A
7 $     24,395 M 30 5 A
8 $     24,395 F 52 6 A
9 $     43,124 M 48 8 A
10 $     23,975 F 58 4 A
11 $     53,174 M 46 4 C
12 $     58,515 M 36 8 C
13 $     56,294 M 49 10 B
14 $     49,033 F 55 10 B
15 $     44,884 M 41 1 A
16 $     53,429 F 52 5 B
17 $     46,574 M 57 8 A
18 $     58,968 F 61 10 B
19 $     53,174 M 50 5 A
20 $     53,627 M 47 10 B
21 $     49,033 M 54 5 B
22 $     54,981 M 47 7 A
23 $     62,530 M 50 10 B
24 $     27,525 F 38 3 A
25 $     24,395 M 31 5 A
26 $     56,884 M 47 10 A
27 $     52,111 M 56 5 A
28 $     44,183 F 38 5 B
29 $     24,967 F 55 6 A
30 $     35,423 F 47 4 A
31 $     41,188 F 35 2 B
32 $     27,525 F 35 3 A
33 $     35,018 M 39 1 A
34 $     44,183 M 41 2 A
35 $     35,423 M 44 1 A
36 $     49,033 M 53 8 A
37 $     40,741 M 47 2 A
38 $     49,033 M 42 10 A
39 $     56,294 F 44 6 C
40 $     47,180 F 45 5 C
41 $     46,574 M 56 8 A
42 $     52,722 M 38 8 C
43 $     51,237 M 58 2 B
44 $     53,627 M 52 8 A
45 $     53,174 M 54 10 A
46 $     56,294 M 49 10 B
47 $     49,033 F 53 10 B
48 $     49,033 M 43 9 A
49 $     55,549 M 35 8 C
50 $     51,237 M 56 1 C
51 $     35,200 F 38 1 B
52 $     50,175 F 42 5 A
53 $     24,352 F 35 1 A
54 $     27,525 F 40 3 A
55 $     29,606 F 34 4 B
56 $     24,352 F 35 1 A
57 $     47,180 F 45 5 B
58 $     49,033 M 54 10 A
59 $     53,174 M 47 10 A
60 $     53,429 F 45 7 B
61 $     53,627 M 47 10 A
62 $     26,491 F 46 7 A
63 $     42,961 M 36 3 B
64 $     53,174 M 45 5 A
65 $     36,292 M 46 0 A
66 $     37,292 M 47 1 A
67 $     41,188 F 34 3 B
68 $     57,242 F 45 7 C
69 $     53,429 F 44 6 C
70 $     53,174 M 50 10 B
71 $     44,138 F 38 2 B
Answer #1
F M
42 54
52 47
58 47
55 44
52 42
61 30
38 48
38 46
55 36
47 49
35 41
35 57
44 50
45 47
53 54
38 47
42 50
35 31
40 47
34 56
35 39
45 41
45 44
46 53
34 47
45 42
44 56
38 38
58
52
54
49
43
35
56
54
47
47
36
45
46
47
50
mean 43.96429 46.55814
sd 7.805421 6.877508
n 28 43

M1 & M2 = sample means
t = t statistic determined by confidence level
s(M1 - M2) = standard error = √((s2p/n1) + (s2p/n2))

μ1 - μ2 = (M1 - M2) ± ts(M1 - M2)

margin or error = ts(M1 - M2)

Calculation

Pooled Variance
s2p = ((df1)(s21) + (df2)(s22)) / (df1 + df2) = 3631.57 / 69 = 52.63

Standard Error
s(M1 - M2) = √((s2p/n1) + (s2p/n2)) = √((52.63/28) + (52.63/43)) = 1.76

Margin of error

= 1.67 * 1.76 = 2.9392

Confidence Interval
μ1 - μ2 = (M1 - M2) ± ts(M1 - M2) = 2.5939 ± (1.67 * 1.76) = 2.5939 ± 2.9372187